# KinesCeTI: A Modular and Size-Adaptable Force Feedback Glove with Interchangeable Actuation for the Index and Thumb

### v1

@@ 6 Which condition do you prefer for this type of task: with the free hand or with the exoskeleton?
Scales: 0–100 with anchors specific to each question. Example anchors:
Q1: 0 = Not, 50 = Moderate, 100 = Extremely helpful.
@@
Q3: 0 = None, 50 = Moderate, 100 = Extreme distraction.
Q4: 0 = None, 50 = Moderate, 100 = Extreme annoyance.
Q5: 0 = Not, 50 = Moderate, 100 = Extremely helpful.
@@
Intermediate anchors (25, 75) were also shown to participants.
@@ IV-B2 Experiment results
The study included 20 participants (16 males, 4 females), aged 26–62 years (mean = 32.25), of whom 18 were right-handed and 2 left-handed. For completion times, mean values were computed per participant for each combination of tolerance and goal. As data were not normally distributed, Wilcoxon signed-rank tests were applied under the alternative hypothesis that the free-hand condition would yield longer times. Bonferroni–Holm correction was applied across the four comparisons. As summarized in Table X, all differences remained statistically significant after correction (p=.0002p=.0002–.0068). The free-hand condition consistently resulted in longer completion times ($\tilde{\Delta}\approx$ +0.4–0.6 s), with large effect sizes in all cases.
Across all trials, more failures (defined as dropping or over-pinching the virtual balls) occurred with the free hand (320) than with the exoskeleton (29). As failures per configuration (goal and threshold) were sparse and commonly zero for the exoskeleton, they were aggregated per participant across all trials and compared. The free-hand condition produced substantially more errors (Mdn = 9.5, IQR = 2.0–31.5) than the exoskeleton (Mdn = 1.0, IQR = 0.0–2.25). A one-sided Wilcoxon signed-rank test confirmed this difference, $W=145$, $z=3.29$, $p<.001$, with a large effect size ($r=.79$).
When comparing data between the two tolerance thresholds (0.025 vs. 0.030), no consistent differences were found. Occasional effects were observed in failures, but these were based on sparse and zero-inflated data and are not considered conclusive. Similarly, no statistically significant differences were observed between Goal 1 and Goal 2 after applying Wilcoxon tests with Bonferroni–Holm correction.
TABLE X: Wilcoxon results for completion times (Leap Motion vs Exoskeleton)
@@
NASA-TLX scores were compared between free-hand and exoskeleton conditions using one-sided Wilcoxon signed-rank tests, under the alternative hypothesis that workload would be higher with the free hand. Results are summarized in Fig. 15 and Table XI. After Bonferroni-Holm correction, significant differences were observed in Mental Demand ($p=.008$), Performance ($p=.02$), Effort ($p=.01$), Frustration ($p=.02$) and the overall NASA-TLX Total score ($p=.01$), all favoring the exoskeleton. Physical Demand and Temporal Demand did not reach significance ($p=.155$ each). These results suggest that the exoskeleton condition was perceived as less demanding overall, with lower mental effort, frustration and effort ratings, as well as better perceived performance.
Figure 15: NASA-TLX results for the virtual pick and place task.
TABLE XI: NASA-TLX scores of virtual pick and place test reported as median [IQR] and Wilcoxon signed-rank results (n=20).
@@
With regards to the custom questionnaire, participants rated force feedback as helpful (Q1: Mdn = 64.5, IQR = 40.0–75.0)). Sounds were clearly noticeable (Q2: Mdn = 75.0, IQR = 25.8–76.3), but rated low in distraction (Q3: Mdn = 5.5, IQR = 1.5–23.5) and annoyance (Q4: Mdn = 5.5, IQR = 2.8–12.8). The perceived helpfulness of sounds varied considerably across participants (Q5: Mdn = 27.5, IQR = 7.8–70.3). Finally, overall preference strongly favored the exoskeleton (Q6: Mdn = 85.0, IQR = 60.8–97.8), well above the neutral midpoint of 50.
These results suggest that the exoskeleton with force feedback enhanced task performance relative to the free-hand condition, with shorter completion times and fewer failures. Subjective measures were consistent with this outcome: workload ratings (NASA-TLX) were generally lower with the exoskeleton and participants considered force feedback helpful, while motor sounds were noticeable but not distracting or annoying. Notably, overall preference favored the exoskeleton. Together, these results suggest that this system enhances performance in virtual pick-and-place tasks and is well accepted by users.
IV-C Softness discrimination study
@@ IV-C1 Experiment definition
This experiment evaluated the one-way clutch mechanism in a pinching softness discrimination task in VR. Force feedback was rendered by a pseudo-impedance controller that engages the clutch to provide unidirectional active forces on the fingers. The control algorithm running on the ESP32 computed the motor command as
$u_{i}\;=\;\theta_{\mathrm{eq}}\;+\;0.01\,k\,s_{i},$ (2)
where $u_{i}$ is the motor command (servo position), $s_{i}$ is the virtual fingertip penetration into the object, $k$ is the softness parameter (pilot-calibrated), $0.01$ is a fixed gain and $\theta_{\mathrm{eq}}$ is the equilibrium motor position when the clutch engages the main pulley after feedback is activated.
Three softness levels were tested, selected via preliminary calibration: rigid (A), medium-rigid (B) and soft (C).The experimental hypothesis was that the one-way clutch system would render perceptually distinct softness levels, allowing participants to discriminate between them with accuracy above chance. The task consisted of presenting participants with two virtual balls per trial, each with a different softness, as shown in Fig. 14(b). After probing both, participants indicated which one felt softer. All pairwise comparisons (AB, AC, BC) were tested, each repeated three times, yielding nine samples per participant. Ball positions (left/right) and stimulus order were randomized. Prior to the task, participants performed a brief training (2 min) to become familiar with the procedure.
Data collection included both objective and subjective measures. Objective data consisted of discrimination accuracy (number of correct and incorrect softer-ball selections). Subjective data included a short questionnaire on task perception, force feedback and acoustic awareness (Table XII).
TABLE XII: Customized Questionnaire for the Pinching Softness Discrimination Task
@@ 7 To what extent were the sounds from the exoskeleton motors helpful during the task?
Scales: 0–100, anchors specific to each question.
Q1: 0 = None, 50 = Moderate, 100 = Extremely clear.
@@
Q3: 0 = Never, 50 = Sometimes, 100 = Always same.
@@
Q5: 0 = None, 50 = Moderate, 100 = Extremely distracting.
Q6: 0 = None, 50 = Moderate, 100 = Extremely annoying.
Q7: 0 = Not helpful, 50 = Moderate, 100 = Extremely.
Intermediate anchors (25, 75) were also provided.
@@ IV-C2 Experiment results
The study included 20 participants (18 males, 2 females), aged 24–39 years (mean = 30.15), of whom 14 were right-handed and 6 left-handed. Performance was assessed as correct or incorrect identification of the softer object. Across all trials, the overall discrimination accuracy was high (Mdn = 88.9%, IQR = 77.8–100.0). As the data were not normally distributed, nonparametric analyses were applied. Because responses were binary (success vs. failure), one-sided binomial tests were conducted against the null hypothesis of chance performance (50%), with Bonferroni-Holm correction across the three pairwise comparisons (AB, AC, BC). Results are summarized in Table XIII. Accuracy was statistically significant above chance for all comparisons (all $p<.001$), with large to very large effect sizes (Cohen’s h=0.69h=0.690.990.99), indicating that participants reliably discriminated between the tested softness levels.
TABLE XIII: Binomial test results for softness comparisons
Comparison Correct (k/n) Proportion 95% CI $p$ (Holm) Cohen’s $h$
@@
In the custom questionnaire, participants generally reported that stiffness differences between the two balls were clearly perceivable (Q1: Mdn = 67, IQR = 48.3–75.0) and expressed moderate to high confidence in identifying the softer ball (Q2: Mdn = 70, IQR = 47.5–85.0). Still, indistinguishability was occasionally reported (Q3: Mdn = 33.5, IQR = 24.8–54.8), indicating that some comparisons remained challenging. Motor sounds were rated as highly noticeable (Q4: Mdn = 75.0, IQR = 58.3–90.3), but distraction (Q5: Mdn = 10.0, IQR = 0.0–18.5) and annoyance (Q6: Mdn = 14.0, IQR = 0.0–23.5) were minimal. Finally, the perceived helpfulness of sounds varied substantially (Q7: Mdn = 25.0, IQR = 9.3–52.0), suggesting that while audible cues were sometimes useful, their contribution was not consistent across participants.
@@
Overall, the results of this study confirm that the exoskeleton with its one-way clutch mechanism rendered perceptually distinct softness levels, that participants discriminated reliably with relatively high accuracy. Subjective ratings indicated moderately high perception of stiffness differences and confidence in judgments, while acoustic feedback was noticeable but not distracting or annoying. Taken together, these findings support the hypothesis that the system provides effective unidirectional force feedback for softness rendering.
V Conclusions
This work presents KinesCeTI, a modular force feedback glove with flexibility in both its thimble attachment system and its actuation capabilities. The glove integrates a bidirectional tendon routing system and a structure adaptable to different hands, enabling the implementation of external actuation modules. Two mechanisms were developed: a ratchet–pawl brake and a novel one-way clutch, first introduced in this work. The system was evaluated in three user studies, confirming ergonomics across hand sizes, functional operation of both mechanisms, the ability to manipulate real and virtual objects and the feasibility of rendering variable impedance.
Compared to existing systems, KinesCeTI offers advantages in various aspects. First, in hand adaptability, through an adjustable linkage structure and strap thimble design, though fingertip thimbles still required different sizes, which could be simplified with an improved strap-based design. Second, in versatility, as the architecture supports multiple force feedback modalities (passive, active, variable, bidirectional), unlike existing solutions such as HaptX and SenseGlove. The type of stimulus is therefore dependent on the externally attached actuator module. In this regard, latency remains an area for improvement, with mean values of 58.5 ms for the ratchet–pawl and 88.7 ms for the clutch, both showing high variability. While the ratchet–pawl mean latency is comparable to or slightly lower than the 62 ms reported for HaptX, the clutch remains higher, particularly when compared to Wolverine (21 ms). Nevertheless, the presented mechanisms are lightweight, compact, affordable and simple. Another distinguishing aspect of this work is the explicit consideration of acoustic perception, addressed both in motor selection and in user studies, a factor rarely considered in prior designs.
The technical tests and user studies also highlighted areas requiring further development: improving force resolution and control, reducing latency and its variability, extending modular and bidirectional actuation to the thumb structure and further optimization of weight, volume and ergonomics. More precise digit-level tracking is also needed to complement the current overall position sensing.
Overall, KinesCeTI provides a versatile platform for haptics research. Its modular design allows different actuators and modalities to be tested within the same platform, avoiding redesigns required by closed architectures. Knuckle Hubs can host alternative actuation modules, while the thimble system allows easy integration of tactile, thermal or multimodal modules. This enables comparative testing of different actuation principles and applications within a single glove.
@@ Acknowledgment
[Uncaptioned image] Pablo Alvarez Romeo received the M.Sc. degree in Electronic Engineering from the University of Zaragoza, Zaragoza, Spain. He is currently a Research Associate with the Chair of Acoustics and Haptics, TU Dresden. His research interests include haptic interfaces and perception, robotics, mechatronics design and development, wearables, additive manufacturing and rapid prototyping.
[Uncaptioned image] Mehmet Ercan Altinsoy received the graduation degree in mechanical engineering from the Technical University of Istanbul, Istanbul, Turkey and the Ph.D. degree in electrical engineering from Ruhr-University Bochum, Bochum, Germany, in 2005. After receiving his Ph.D. degree, he was with HEAD Acoustics as a Consulting Engineer. Since 2006, he has been with TU Dresden, Dresden, Germany, where he is currently a Professor in acoustic and haptic engineering. His research interests include perception-based engineering, vibroacoustics, vehicle acoustics, electroacoustics, haptic interfaces, haptic perception, whole-body vibrations, product sound and vibration design. Dr. Altinsoy was a member of the International Graduate School for Neuroscience with Ruhr-University Bochum. In 2018, he was awarded a Visiting Professorship from Tohoku University, Japan. He is a Lothar-Cremer Medalist of the Acoustical Society of Germany, DEGA. He is the Chairman of the Vehicle NVH Expert Committee of DEGA and one of the core team members of the Cluster of Excellence Centre for Tactile Internet with Human-in-the-Loop.

### v2

@@ 6 Which condition do you prefer for this type of task: with the free hand or with the exoskeleton?
Scales: 0–100. Intermediate anchors (25, 75) were also provided.
Q1, Q5: 0 = Not, 50 = Moderate, 100 = Extremely.
@@
Q3, Q4: 0 = None, 50 = Moderate, 100 = Extreme.
@@ IV-B2 Experiment results
The study included 20 participants (16 males, 4 females), aged 26–62 years (mean = 32.25), of whom 18 were right-handed and two left-handed. For completion times, mean values were computed per participant for each combination of tolerance and goal. As data were not normally distributed, Wilcoxon signed-rank tests were applied under the alternative hypothesis that the free hand condition would yield longer times. Bonferroni–Holm correction was applied across the four comparisons. As summarized in Table VIII, all differences remained statistically significant after correction ($p$ range: .0002.0068.0002–.0068). The free hand condition consistently resulted in longer completion times ($\tilde{\Delta}\approx$ +0.4–0.6 s), with large effect sizes in all cases.
Across all trials, more failures occurred with the free hand (320) than with the exoskeleton (29). As failures per configuration (goal and threshold) were sparse and commonly zero for the exoskeleton, they were aggregated per participant across all trials. The free hand condition produced substantially more failures (Mdn = 9.5, IQR = 2.0–31.5) than the exoskeleton (Mdn = 1.0, IQR = 0.0–2.25). A one-sided Wilcoxon signed-rank test confirmed this difference, $W=145$, $z=3.29$, $p<.001$, with a large effect size ($r=.79$).
When comparing data between the two tolerance thresholds (0.025 vs. 0.03), no consistent differences were found. Occasional effects were observed in failures, but these were based on sparse and zero-inflated data and are not considered conclusive. Similarly, no statistically significant differences were observed between Goal 1 and Goal 2 after applying Wilcoxon tests with Bonferroni–Holm correction.
TABLE VIII: Wilcoxon results for completion times (Leap Motion vs Exoskeleton)
@@
NASA-TLX scores were compared between free hand and exoskeleton conditions using one-sided Wilcoxon signed-rank tests, under the alternative hypothesis that workload would be higher with the free hand. Results are summarized in Fig. 14 and Table IX. After Bonferroni-Holm correction, significant differences were observed in Mental Demand ($p=.008$), Performance ($p=.02$), Effort ($p=.01$), Frustration ($p=.02$) and the overall NASA-TLX Total score ($p=.01$), all favoring the exoskeleton. Physical Demand and Temporal Demand did not reach significance ($p=.155$ each). These results suggest that the exoskeleton condition was perceived as less demanding overall, with lower mental demand, frustration, and effort ratings, as well as better perceived performance.
Figure 14: NASA-TLX results for the virtual pick-and-place task.
TABLE IX: NASA-TLX scores of virtual pick-and-place test reported as median [IQR] and Wilcoxon signed-rank results (n=20).
@@
In the custom questionnaire, participants rated force feedback as helpful (Q1: Mdn = 64.5, IQR = 40.0–75.0). Sounds were clearly noticeable (Q2: Mdn = 75.0, IQR = 25.8–76.3), but rated low in distraction (Q3: Mdn = 5.5, IQR = 1.5–23.5) and annoyance (Q4: Mdn = 5.5, IQR = 2.8–12.8). The perceived helpfulness of sounds varied considerably across participants (Q5: Mdn = 27.5, IQR = 7.8–70.3). Finally, overall preference strongly favored the exoskeleton (Q6: Mdn = 85.0, IQR = 60.8–97.8), well above the neutral midpoint of 50.
These results indicate that the force feedback condition improved task performance compared to the free hand condition, yielding shorter completion times and fewer failures. Subjective measures aligned with these findings: workload (NASA-TLX) was generally lower, force feedback was perceived as helpful, and motor sounds were noticeable but not distracting. Overall user preference favored the exoskeleton, supporting its effectiveness and acceptance in virtual pick-and-place tasks.
IV-C Perceived softness discrimination study
@@ IV-C1 Experiment definition
This experiment evaluated the one-way clutch mechanism in a pinching softness discrimination task in VR. Participants were instructed to probe two virtual spheres per stimulus and indicate which one felt softer (Fig. 13(b)). In this study, softness refers to the perceived compliance of the virtual object, arising from different position-based motion resistance profiles rendered by the exoskeleton, rather than from physically calibrated stiffness values. The fingertip position in CHAI3D was updated using AS5600 angle measurements from the physical exoskeleton. Upon contact, CHAI3D sent an activation signal and the fingertip penetration depth to the ESP32. This value was used in a closed-loop position control algorithm running on the ESP32 to generate unidirectional active resistance on the digits. The position command for each servomotor is computed as
$\theta_{i,\mathrm{cmd}}\;=\;\theta_{\mathrm{eq}}\;+\;0.01\,c\,s_{i},$ (2)
where $\theta_{i,\mathrm{cmd}}$ is the servomotor position command expressed in internal control units (0–600 uds), corresponding to a physical motion range of 0 to $120^{\circ}$. The variable $s_{i}$ is the virtual fingertip indentation into the object expressed in simulation units defined by the virtual object geometry and hand model (1: surface contact, 295: full pinch closure), $c$ is an empirically calibrated compliance parameter, $0.01$ is a fixed gain, and $\theta_{\mathrm{eq}}$ is the equilibrium motor position when the clutch completes engagement and feedback is activated. Three compliance levels were tested, selected via preliminary calibration: rigid (A, $c=10$), medium-rigid (B, $c=50$) and soft (C, $c=90$). Lower values of $c$ restrict indentation and emulate higher stiffness, whereas higher values of $c$ allow deeper indentation.
The hypothesis was that the system would render perceptually distinct softness levels, enabling discrimination above chance ( 50%). All pairwise comparisons (AB, AC, BC) were tested with three repetitions each, yielding nine samples per participant, with randomized stimulus order and left/right placement. Prior to the task, participants performed a brief training (2 min) to become familiar with the procedure. Data collection included both objective and subjective measures. Objective data consisted of discrimination accuracy, defined as the proportion of correct softer-ball selections per condition. Subjective data included a short questionnaire on task perception, force feedback and acoustic awareness (Table X).
TABLE X: Questionnaire (Study 3)
@@ 7 To what extent were the sounds from the exoskeleton motors helpful during the task?
Scales: 0–100. Intermediate anchors (25, 75) were also provided.
Q1: 0 = No difference, 50 = Moderate differences, 100 = Extremely clear.
@@
Q3: 0 = Never, 50 = Sometimes, 100 = Always.
@@
Q5-7: 0 = Not at all, 50 = Moderate, 100 = Extremely.
@@ IV-C2 Experiment results
The study included 20 participants (18 males, 2 females), aged 24–39 years (mean = 30.15), of whom 14 were right-handed and 6 left-handed. Performance was assessed as correct or incorrect identification of the softer object. Across all trials, the overall discrimination accuracy was high (Mdn = 88.9%, IQR = 77.8–100.0). As the data were not normally distributed, nonparametric analyses were applied. Because responses were binary (success vs. failure), one-sided binomial tests were conducted against the null hypothesis of chance performance (50%), with Bonferroni-Holm correction across the three pairwise comparisons (AB, AC, BC). Results are summarized in Table XI. Accuracy was statistically significant above chance for all comparisons (all $p<.001$), with large to very large effect sizes (Cohen’s h=0.69h=0.690.990.99), indicating that participants reliably discriminated between the tested softness levels.
TABLE XI: Binomial test results for softness comparisons
Pair Correct Proportion Clopper–Pearson 95% CI $p$ (Holm) Cohen’s $h$
@@
In the custom questionnaire, participants generally reported that stiffness differences between the two balls were clearly perceptible (Q1: Mdn = 67, IQR = 48.3–75.0) and expressed moderate to high confidence in identifying the softer ball (Q2: Mdn = 70, IQR = 47.5–85.0). Still, indistinguishability was occasionally reported (Q3: Mdn = 33.5, IQR = 24.8–54.8), indicating that some comparisons remained challenging. Motor sounds were rated as highly noticeable (Q4: Mdn = 75.0, IQR = 58.3–90.3), but distraction (Q5: Mdn = 10.0, IQR = 0.0–18.5) and annoyance (Q6: Mdn = 14.0, IQR = 0.0–23.5) were minimal. Finally, the perceived helpfulness of sounds varied substantially (Q7: Mdn = 25.0, IQR = 9.3–52.0), suggesting that while audible cues were sometimes useful, their contribution was not consistent across participants.
@@
Overall, results indicate that the one-way clutch mechanism rendered perceptually distinct compliance levels that participants reliably discriminated in pairwise comparisons. Subjective ratings indicated moderately high perception of stiffness differences and confidence in judgments, while acoustic feedback was noticeable but not distracting or annoying. Taken together, these findings support the hypothesis that the proposed system can provide perceptually meaningful unidirectional force feedback for softness discrimination. As no objective interaction measures were recorded, conclusions are limited to subjective perception and relative discriminability.
V Discussion
The experimental evaluations present positive results regarding the intended design goals and align with prior haptic glove studies. Ergonomics, a key focus in commercial systems such as SenseGlove and HaptX, were evaluated favorably, demonstrating adaptation to different hand sizes and largely unrestricted motion across tested gestures at both fingertip and second-phalanx attachment points. However, larger and more gender-balanced studies are required to assess broader applicability. User studies confirmed perceptually meaningful force feedback for both actuation mechanisms. The ratchet–pawl mechanism provides rigid motion constraints similar in principle to those in prior systems such as Dexmo and Wolverine, although stiffness was not quantitatively evaluated, while the third study showed that the clutch mechanism can render perceptually distinct compliance levels. In the pick-and-place task, failure rates were lower with force feedback compared to the free-hand condition, indicating improved grasp control. This is consistent with prior work on Dexmo [19] in VR and Leonardis et al. [27] in teleoperation, where task success or error rates improved with force feedback. The latter also reported reduced execution times despite differences in the experimental setup.
At the system level, commercial devices such as HaptX G1 and SenseGlove Nova 2 have comparable weight to KinesCeTI (450 g and 350 g vs. 376 g) and lower latency (HaptX G1: 62 ms vs. 83.15 ms ratchet–pawl, 100.1 ms clutch) but offer limited force feedback dimensionality. Linkage-based systems in teleoperation and rehabilitation [27, 6, 22] provide effective force feedback but rely on external, immovable actuation, limiting portability, whereas KinesCeTI supports multiple feedback modalities in a portable, wearable architecture.
Adaptability to different users has been addressed in prior work through adjustable components [10], multiple glove sizes (HaptX, Senseglove), or size-adaptable exoskeleton designs [22]. The proposed linkage structure extends this adaptability by supporting attachment at both the fingertip and the second phalanx, whereas prior designs focus on a single contact location. Together, this dual-point adaptability and the consideration of anthropometric extremes introduce a key trade-off: increased volume. This increases the likelihood of collisions in constrained environments and contributes to mass and inertia that are perceptible to users, particularly due to extensive use of bearings and prototyped electronics. These effects could be mitigated through structural optimization and compact PCB-based electronics, which would also free dorsal space and enable extension to additional digits. The thumb assembly, despite incorporating a compression spring to reduce inertial effects, would particularly benefit from such refinements. As a grounded exoskeleton, the system transmits reaction forces to the hand, which remained noticeable despite a compliant TPU–neoprene attachment and could be further reduced through optimizations. Another limitation is the absence of integrated force sensing, necessary for closed-loop force control. The main pulley offers a natural location for integrating force sensing within future actuation modules.
Despite these limitations, KinesCeTI adopts a different approach from many existing solutions that rely on closed architectures with fixed capabilities or prioritize specific features at the expense of adaptability and wearability. Its contribution therefore lies not in outperforming existing systems in isolated metrics, but in providing a versatile and modular platform that enables researchers to integrate and evaluate novel haptic actuation and feedback concepts, thereby lowering the entry barrier for haptic research. Further contributions of this work include the actuation mechanisms, the explicit consideration of acoustic perception, and the evaluation protocols, all of which may be reused or adapted for other hand-based haptic interfaces, potentially extending the impact of this work.
VI Conclusions
This work presents KinesCeTI, a modular and portable force feedback glove designed as a flexible research platform. It integrates an adaptive linkage structure, bidirectional tendon routing, and interchangeable thimble and actuation modules, enabling the exploration of multiple force feedback modalities. Two mechanisms were developed: a ratchet–pawl brake and a novel one-way clutch, first presented in this work.
The system was evaluated in three user studies addressing ergonomics, hand motion, and interaction performance in real and virtual environments. The results demonstrated that KinesCeTI can be worn by users with different hand sizes while being attached at either the fingertip or the second phalanx, and that both actuation modules rendered perceptually meaningful force feedback during manipulation tasks.
Rather than targeting peak performance in individual metrics, KinesCeTI is conceived as an open and adaptable platform for haptics and VR research. Its modular architecture enables comparative testing of multiple actuation concepts within a single glove, either at the Knuckle Hubs or through interchangeable thimbles that allow integration of tactile, thermal, or multimodal feedback modules without structural redesign. This reduces the effort required to develop and systematically evaluate diverse haptic systems.